1. If you mix 10 mL of 0.020 M AgNO3 and 10mL of 0.020 M NaCl, which statement is true?

a. Q > K, ppt forms

b. Q < K, ppt forms

c. Q > K, ppt does not form

d. Q < K, ppt does not form

e. Q is unknowable
Answer #1

Since Volume doubles on mixing, concentration becomes half.

[Ag] = [Cl] = initial molarity / 2 = 0.01 M

Q = [Ag+] * [Cl-] = 0.01 * 0.01 = 10-4 M2

Ksp = 1.77 * 10-10 M2

Therefore, Q > K and ppt forms